# MetaphorShare: A Dynamic Collaborative Repository of Open Metaphor Datasets

### v2

@@ 5 Evaluation
Then, as a case study we perform a cross-dataset metaphor identification analysis to illustrate a possible usage of MetaphorShare for NLP research. Given an input expression, the task consists of deciding whether the expression is a metaphor or not. Different projects label metaphors for different purposes. MetaphorShare makes it easy to fine-tune models on relevant dataset to further support project-specific automatic or semi-automatic labeling.
@@
In order to maintain a training set of the same size for all datasets, we randomly sampled 800 examples from each set to create the training sets. Because our datasets from psycholinguisics are small and similar in their creation method, we grouped them into one single set called J&C in this experiment. RoBERTa base models Liu et al. (2019) are then finetuned independently on the 10 datasets and tested on each of them. Similarly to the experiments in Boisson et al. (2023), for hyperparameter optimisation, we rely on the Bayesian Optimization with the Hyperband (BOHB) algorithm Falkner et al. (2018) available in RayTune Liaw et al. (2018), with an identical hyperparameter search space and 25 trials.
@@ Results.
Results are shown in Table 4. As expected, models trained and evaluated on the same dataset often achieved the best results. A few datasets generalise better than others. It is surprisingly the case for J&C and GUT, both containing only short example with constrained syntactic structures. On the other hand, datasets such as TONG and NEWS do not generalize as well, probably due to different labelled PoS, to the token or MWE span of the labels, or due to redundant contexts corresponding to datasets created for other tasks such as paraphrasing in the case of TONG.
@@ 6 Conclusion and Future Work
In this paper, we have presented a website to label, unify and share metaphor datasets. It enables an easy integration of new resources of different original formats by the community. The 25 currently included datasets are in English and we are planning to integrate and reach communities working on other languages in the future as one of the main aims.
In the future we will focus on integrating a functionality for automatically labelling raw text. The progress recently made by transformer-based language models for metaphor identification opens the path to the creation of helpful personalized tools for automatic or semi-automatic labelling of metaphors based on an initial sample of annotations.
@@ Acknowledgments
This research was supported by the Association for Researching and Applying Metaphors through the Building Bridges fund. We thank the British National Corpus for giving us the permission to share fragments of texts of the BNC directly on MetaphorShare. We also thank Eileen Cardillo, all the authors of Gutiérrez et al. (2016) and of Birke and Sarkar (2006) for allowing us to redistribute their datasets online for research purposes, and to the rest of the authors of the datasets included in the repository for the creation of metaphor datasets with a flexible open license. Without the effort of the metaphor community, this work would not have been possible.
Appendix A Detailed information of datasets upoaded in MetaphorShare and used in our experiments
Table 5 provides detailed information about the datasets, including license, source corpora, label distribution and sentence distribution. Additional links and preprocessing details are listed below :
@@ Appendix A Detailed information of datasets upoaded in MetaphorShare and used in our experiments
Anomalous sentences are not shown in the table.
@@ Appendix B Dataset upload
Figure 3 & 4 show the row by row feedback provided after the automatic validation step when the file is rejected, and the pop up message indicating that a file passed the automatic validation step.
@@
Additional Information attached to a labelled example after validation.
@@ Appendix D Online annotation tool
Figure 5 shows the annotation tool interface with one example sentence. Figure 6 shows how the created records of a dataset are displayed, allowing an easy step by sep construction of a dataset.
@@
Figure 5: Screenshot of the online annotation tool showing the text input area, tag selection and creation, and coloured resulting tagged text.
[…]

### v3

@@ 5 Evaluation
Then, as a case study we perform a cross-dataset metaphor identification analysis to illustrate a possible usage of MetaphorShare for NLP research. Given an input expression, the task consists of deciding whether the expression is a metaphor or not. Different projects label metaphors for different purposes. MetaphorShare makes it easy to fine-tune models on specific datasets to further support project-specific automatic or semi-automatic labeling.
@@
In order to maintain a training set of the same size for all datasets, we randomly sampled 800 examples from each set to create the training data. Because our datasets from psycholinguisics are small and similar in their creation method, we grouped them into one single set called J&C in this experiment. RoBERTa base models Liu et al. (2019) are then finetuned independently on the 10 datasets and tested on each of them. Similarly to the experiments in Boisson et al. (2023), for hyperparameter optimisation, we rely on the Bayesian Optimization with the Hyperband (BOHB) algorithm Falkner et al. (2018) available in RayTune Liaw et al. (2018), with an identical hyperparameter search space and 25 trials.
@@ Results.
Results are shown in Figure 3. As expected, models trained and evaluated on the same dataset often achieved the best results. A few datasets generalise better than others. It is surprisingly the case for J&C and GUT, both containing only short example with constrained syntactic structures. On the other hand, datasets such as TONG and NEWS do not generalize as well, probably due to different labelled PoS, to the token or MWE span of the labels, or due to redundant contexts corresponding to datasets created for other tasks such as paraphrasing in the case of TONG.
@@ 6 Conclusion and Future Work
In this paper, we have presented a website to label, unify and share metaphor datasets. It enables an easy integration of new resources of different original formats by the community. There are currently 25 English datasets integrated in the platform. In the future, we are planning to integrate and reach communities working on other languages as one of the main aims.
As far as the platform is concerned, we will focus on integrating a functionality for automatically labelling raw text. The progress recently made by transformer-based language models for metaphor identification opens the path to the creation of helpful personalized tools for automatic or semi-automatic labelling of metaphors based on an initial sample of annotations.
@@ Acknowledgments
This research was supported by the Association for Researching and Applying Metaphors through the Building Bridges fund. We thank the British National Corpus for giving us the permission to share fragments of texts of the BNC directly on MetaphorShare. We also thank Eileen Cardillo, all the authors of Gutiérrez et al. (2016) and of Birke and Sarkar (2006) for allowing us to redistribute their datasets online for research purposes, and to the rest of the authors of the datasets included in the repository for the creation of metaphor datasets with a flexible open license. Without the effort of the metaphor community, this work would not have been possible. Jose Camacho-Collados was supported by a UKRI Future Leaders Fellowship.
Appendix A Detailed information of datasets used in our experiments
Table 4 provides detailed information about the datasets, including license, source corpora, label and sentence distribution. Additional links and preprocessing details are listed below :
@@ Appendix A Detailed information of datasets used in our experiments
Anomalous sentences and simile are not shown in the table. because because they are not used in our binary classification evaluation task.
@@ Appendix B Dataset upload
Figure 5 & 6 show the row by row feedback provided after the automatic validation step when the file is rejected, and the pop up message indicating that a file passed the automatic validation step.
@@ Appendix D Online annotation tool
Figure 7 shows how the created records of a dataset are displayed, allowing an easy step by step annotation.
Dataset Reference License N N dist. N dist. % Expr. Domain/Source
Name ctxt. expr. met. PoS
Words in syntactically constrained sentences (Psycholinguistics)
JANK Jankowiak (2020) CC BY 4.0 240 240 120 50 N constructed examples
CARD_V Cardillo et al. (2010) CC BY-NC 280 280 140 50 V constructed examples
CARD_N Cardillo et al. (2010, 2017) CC BY-NC 512 512 256 50 N constructed examples
Words in natural short contexts (NLP)
MOH Mohammad et al. (2016) see data page 1632 1632 439 25 V WordNet examples
NewsMet Joseph et al. (2023) Apache-2.0 1205 1205 477 49 V Fake News Corpus
TSV_A Tsvetkov et al. (2014) see data page 1945 1072 687 50 A various websites
GUT Gutiérrez et al. (2016) AFL-3.0 8591 3479 23 54 A Wikipedia, UKWaC…
Multi-Word Expressions in natural long contexts (NLP)
PVC Tu and Roth (2012) no license 1348 1348 23 65 V-Prep BNC
MAD Tayyar Madabushi et al. (2022b) GPL-3.0 4558 4554 251 48 NC Common Crawl
MAGPIE Haagsma et al. (2020) CC-BY-4.0 48395 47283 9307 75 Various BNC, PMB.
Words sampled from VUAC (MIPVU)
VUAC_BO Boisson et al. (2023) CC BY-SA 3.0 39223 11476 8674 52 Various VUAC
TONG Tong et al. (2024) CC-BY-4.0 1428 739 861 46 Various VUAC & paraphrases
Table 4: Twelve example metaphor datasets included in the MetaphorShare repository. The N column shows the number of instances in the datasets, followed by the number of distinct provided context surrounding the potential metaphoric expression (N dist ctxt), and the number of distinct lemmas labelled metaphoric or literal (N dist. expr.). %met. indicates the percentage of expressions labelled as metaphorical, and Expr. PoS their part of speech.